OpenStudy (ashsmashum1290):

A donut has a circumference of 25.12 what is the donuts radius? use 3.14 A.4 cm B.8 cm C.12 cm D.16 cm

TheSmartOne (thesmartone):

Hint: \(\sf Circumference = 2 \times \pi \times radius\)

OpenStudy (alexalas):

C=2πr C=25.12

OpenStudy (ashsmashum1290):

so 2x3.14 but how do i get the radius

TheSmartOne (thesmartone):

So you see they gave you the circumference. And that is equal to 25.12 So plug it all in to the formula: \(\sf 25.12 = 2 \times 3.14 \times radius\)

TheSmartOne (thesmartone):

And then you just divide to isolate radius by itself and you have your answer :)

OpenStudy (ashsmashum1290):

so what is the radius??????

TheSmartOne (thesmartone):

You tell me.

OpenStudy (alexalas):

its simple 25.12/2=3.14 x radius

TheSmartOne (thesmartone):

Do the division \(\sf \Large\frac{25.12}{2 \times 3.14}\)

OpenStudy (ashsmashum1290):

4

TheSmartOne (thesmartone):

Correct :)

OpenStudy (ashsmashum1290):

thanks you

TheSmartOne (thesmartone):

(: